A gas 'x' is collected over water at 17 degree celcius and 750 mm pressure.If the volume of the gas collected is 50 cc then calculate the volume of the dry gas at standard temperature pressure and at 17 degree celcius the vapour of pressure is 14 mm
1
Expert's answer
2011-10-07T08:23:15-0400
Averagemolecular weight of amino acid (AA) is 105 DaAveragemolecular weight of nucleotide(Nucl.) is 345 Da
Averagelength of nucleotide is 0,34 nm
N(AA) = 40000/105 = 381
N(Nucl.) = 381*3 = 1143
l(Nucl.) = 1143*0.34 = 388.62nm
m(Nucl.)1strand = 1143*345 = 394335Da = 394.3 kDa
m(Nucl.)2strand = 1143*345*2 = 788670Da = 788,7 kDa